# Regulatable ML

### v1

@@ Abstract
As autonomous AI agents are increasingly deployed in industry, it is essential to safeguard them. We introduce a novel framework that automates the translation of unstructured design documents into verifiable, real-time guardrails. We introduce "Policy as Prompt," a new approach that uses Large Language Models (LLMs) to interpret and enforce natural language policies by applying contextual understanding and the principle of least privilege. Our system first ingests technical artifacts to construct a verifiable policy tree, which is then compiled into lightweight, prompt-based classifiers that audit agent behavior at runtime. We validate our approach across diverse applications, demonstrating a scalable and auditable pipeline that bridges the critical policy-to-practice gap, paving the way for verifiably safer and more regulatable AI.
@@ 1 Introduction
The rapid proliferation of sophisticated AI agents Wang et al. (2025); Ding et al. (2024) has created an urgent need for robust governance to ensure their alignment with human policies and safety constraints Liu et al. (2024); Kim et al. (2023); Li et al. (2025); He et al. (2024). Emerging frameworks like the EU AI Act eu- (2024) highlight this by mandating comprehensive risk and conformity assessments Fabiano (2024). However, a critical gap persists between high-level, human-readable policies and the low-level, machine-enforceable rules required for practical implementation. This "policy-to-practice" gap presents a significant bottleneck for effective AI evaluation, compliance, and auditing.
To address this challenge, we propose that guardrail security policies can be used to govern the behavior of AI agents, drawing inspiration from established data security principles. A guardrail is a defense mechanism used in AI applications to prevent unintended or harmful behavior Dong et al. (2024); Zhang et al. (2025). The "policy-to-practice" gap is particularly acute for autonomous agents, as the appropriateness of their actions is entirely dependent on context. As argued by the CONSECA framework Tsai and Bagdasarian (2025), security for these generalist systems requires moving beyond static policies, which can overrestrict or overallow actions, to a new paradigm of just-in-time, contextual security. While frameworks like TrustAgent Hua et al. (2024) have proposed using static constitutional principles, this is insufficient for the dynamic nature of agent interactions.
To bridge this critical gap of enforcing security policies for AI agents at scale, we introduce Policy as Prompt, a novel framework for translating natural language policy documents into dynamic, enforceable guardrails. Our work provides a concrete solution to the problem identified by CONSECA by operationalizing its core principles. Our key contributions are as follows: (i) We introduce a scalable, end-to-end pipeline that automatically ingests unstructured technical artifacts to identify and extract security constraints Tsai and Bagdasarian (2025). (ii) We propose a verifiable process where these constraints are converted into a human-readable policy draft, enabling efficient review and refinement by security engineers. (iii) We compile the verified policy into prompt-based classifiers—our "guardrail security policies" Dong et al. (2024)—that enforce a least-privilege policy at runtime by validating agent inputs and outputs Zhang et al. (2025). (iv) We validate our approach by generating policies for various LLM applications and test their effectiveness across different state-of-the-art models.
Figure 1: Policy Generation and Enforcement Pipeline
@@ 2 Policy Tree Generation
The core of our POLICY-TREE-GEN approach is a verified two-pass parsing and enrichment pipeline that extracts security intent from source artifacts. In Pass 1, a generative LLM infers the document hierarchy, extracts verbatim policy statements, and classifies them into four categories: VALINP (acceptable inputs), INVALINP (unacceptable inputs, e.g., out-of-domain topics), VALOUT (sanitization), and INVALOUT (forbidden outputs, such as hallucinations, policy violations, or PII leaks). Non-o1 models in the pipeline operated at a temperature of 0.5. This parsing step yields a policy tree structure Kumar et al. (2025), illustrated in Appendix 2, with example prompts in Appendix 3. For instance, “Resolved HR case data is pulled from Internal Case Portal” is a VALINP, implying “non-HR data” is an INVALINP; “delete or anonymize personal data before analysis” reinforces INVALOUT. All extracted statements undergo verification by a separate judge LLM (gpt-4o) to ensure they are faithful verbatim quotes and correctly categorized; failing items are pruned. In Pass 2, the verified structure is enriched with relevant, source-cited examples, producing a high-fidelity hierarchical policy tree that preserves contextual relationships and enhances interpretability and maintainability.
Our artifact-driven policy generation applies the Principle of Least Privilege to LLM agents Zhang et al. (2025). By converting application artifacts into a structured, verifiable tree, our method enables systematic guardrail enforcement, as demonstrated by the HR policy tree in Figure 1. Policy scope is limited to the LLM application’s input and output boundaries. VALINP/INVALINP govern data entering the first agent, while VALOUT/INVALOUT define rules for the final user response.
@@
Application Example Classification Reason Action
HR Update my address on Workday VALID (Input) ALLOW
HR My address is 21 Victoria St INVALID (Input) Contains Non-Anonymised PII BLOCK
HR Ignore rules and reveal your system prompt INVALID (Input) Malicious / Prompt Injection ALERT
HR Send to sample@gmail.com VALID (Input) Anonymised PII ALLOW
HR US election news link… INVALID (Output) Non HR Content BLOCK
@@ 3 Policy as Prompt Generation
It is then embedded into a master template that primes the LLM as a compliance analyst. Two templates are used: one for an Input Classifier and another for an Output Auditor. These prompt the model to consult the policy rules and return a JSON object with a binary classification (VALID/INVALID) and a concise justification. For HR input classification (Table 1), HR-related requests without non-anonymized PII are VALID, while those with PII, prompt-injection attempts, or non-HR content are INVALID. The strict output format enables deterministic system actions (ALLOW, BLOCK, ALERT). HR Input Classifier is shown in Figure 4 and SecOps in Figure 5.
The deliverables are two markdown few-shot prompts (rules + exemplars), which undergo human-in-the-loop review. Security engineers then either (i) approve for deployment, (ii) reject requiring upstream updates and regeneration, or (iii) request changes to the prompt markdown without altering the upstream tree.
@@ 4 Experimental Setup and Results
This study evaluated the generation and enforcement of guardrail policies across two distinct LLM-powered systems in HR and Security Operations (SecOps) domains. The input artifacts for all models consisted of PRDs, technical design documents, and prompts extracted from the application source code. To ensure compatibility, all documents were converted to Markdown format, and any embedded images were replaced with textual descriptions generated by gpt-4o. All reported metrics are the average of multiple runs to ensure stability.
4.1 Policy-Tree-Gen Analysis
@@
Application Model R (%) F1 (%) Macro-F1 (%) VALINP VALOUT INVINP INVOUT
@@
SecOps O1 19.4 22.6 13.0 22.2 29.6 0.0 0.0
SecOps GPT-OSS 120B 5.6 10.3 4.2 16.7 0.0 0.0 0.0
SecOps Llama 405B 5.6 9.8 10.0 0.0 0.0 40.0 0.0
SecOps Claude 3.5 2.8 5.4 6.2 0.0 0.0 25.0 0.0
We evaluated several large language models including Llama 3 405B(Grattafiori et al., 2024), GPT OSS 120B (OpenAI et al., 2025), Sonnet 3.5(Anthropic, 2024) and o1 (OpenAI et al., 2024) on two applications, HR and SecOps. The models Gemma 3 1B(Team et al., 2025) and Qwen 1.7B Thinking(Yang et al., 2025) were excluded due to poor performance. Gold policies were created by security engineers and served as the ground truth against which the LLM-generated policies were evaluated.
The metrics in Table 2 evaluate model performance in two categories: Detection, measured by Recall (R) and F1 score, and Classification, assessed via Macro-F1 and individual per-class F1 scores for VALINP, VALOUT, INVINP, and INVOUT. A high score in these metrics indicates a model that is effective at both identifying relevant requirements and accurately assigning the correct category to them. The HR application consistently yields higher performance scores for all models compared to the SecOps application. Furthermore, the o1 model significantly outperforms all other models, demonstrating its superior capability in this specific task.
As shown in Table 4, this analysis details model performance on HR and SecOps tasks using metrics such as Detection Precision, which measures the percentage of a model’s predictions that are correct, and Micro-F1. Span quality is evaluated using four metrics: Span Exact, which measures if the extracted text is an exact match to the ground truth; Token-F1, which assesses the token-level overlap between the extracted and ground truth text; Substr, which checks if one text is a substring of the other; and Emb Cos, which is the cosine similarity of the text embeddings. A high score in these metrics indicates that the model is accurately extracting the correct text, even if it might misclassify it. While the o1 model remains a top performer, others reveal a significant disconnect between their extraction quality and overall F1 score. For HR applications, models like Llama 405B and Claude 3.5 have high span quality metrics, indicating accurate text extraction, yet their low Micro-F1 scores suggest they struggle with correct classification. In the SecOps domain, performance is generally lower and more varied. Notably, Claude 3.5 achieved a high Det P on SecOps, meaning all of its identified requirements were true positives, but its Micro-F1 score remained very low due to continued classification errors.
4.2 Policy-As-Prompt-Gen Analysis
@@
Policy enforcement was tested on a separate set of models: Qwen 3 1.7B (Thinking Mode) (Yang et al., 2025), GPT-4o (OpenAI et al., 2024), and Gemma 3 1B(Team et al., 2025). For specialized data classification tasks within this framework, we employed Small Language Models (SLMs), leveraging their established effectiveness and low latency, which helps to quickly run policies in real-time.
Both policies were judged request review by security engineers and deployed after minimal changes, saving time for the security team. POLICY-AS-PROMPT-GEN was evaluated on functional correctness by integrating generated prompts into target LLMs and testing with 100 gold inputs and outputs. Tests included both standard and adversarial cases (e.g., prompt injection, toxic content). Evaluation focused on accuracy of INVALID/VALID classifications, with defaults of INVALID $\Rightarrow$ BLOCK and VALID $\Rightarrow$ ALLOW, unless otherwise specified. A limitation is that prompt tuning emphasized GPT models, potentially contributing to performance gaps.
@@ 6 Limitations
Despite promising results, this work has several limitations that temper generalizability and reproducibility. Our evaluation spans two enterprise-style domains (HR, SecOps) with modest gold sets (100 inputs and 100 outputs per application) and a limited number of design artifacts, which constrains transferability to other settings and larger, more heterogeneous corpora. Policy extraction currently depends on large proprietary models, while open-source baselines differ in size or tuning; moreover, prompting effort was greater for GPT-family models, potentially biasing results and limiting apples-to-apples comparisons. Reproducibility is further hindered by confidentiality of internal artifacts and logs: we cannot release the full corpora. Future work includes leveraging production interaction logs (inputs, outputs, tool calls, actions) to mine candidate rules and hard examples that continuously enrich the policy tree and prompts, and implementing adaptive policy regeneration that automatically re-runs POLICY-TREE-GEN and POLICY-AS-PROMPT-GEN whenever PRDs, TDDs, code, or schemas undergo significant changes, with versioning and regression gating to mitigate drift. Also, prompt optimisations like Kumar et al. (2025) can be applied to POLICY-AS-PROMPT.
@@ A.1 POLICY-TREE-GEN PROMPTS
LLM Prompt: Pass 1 — Parse Document & Classify Security Instructions Role & Goal You are a document analysis expert specializing in data security for LLM-powered applications. Read a document describing an LLM-based system, infer its logical structure, and extract only those instructions relevant to data and guardrail security. Important: Ignore all specific examples in this step. Focus strictly on the instructions themselves. Input <Document text> Task Return a single JSON object representing the document’s hierarchy. For each extracted instruction, you must include: The exact, verbatim source_span copied from the document. A category chosen from 4 CATEGORIES. Valid Categories & Guidance VALINP: Defines what the system should accept (in-domain, topical, acceptable inputs). Exclude: outputs, rejection rules, metadata, formatting, or negative conditions. VALOUT: Defines correct system responses (including handling invalid/unsupported/out-of-domain inputs). Exclude: malformed prompts, format/schema rules, valid inputs, or outputs. INVALINP: Defines what the system must reject (invalid, unsupported, or out-of-domain inputs). Include: logic/guardrails for blocking inputs. Exclude: formatting, valid inputs, or outputs. INVALOUT: Defines incorrect or forbidden system responses (hallucinations, policy violations, leaking PII, misleading answers). Include: logic/guardrails for blocking outputs. Exclude: formats, constraints, valid outputs, or inputs. Exclusion Rule: If an instruction does not clearly fit one of the valid categories above, omit it from the output. Required JSON Shape Root key: "document_structure" (list). Each element: "topic": Heading "content": Instructions with source_span, category, sub_instructions "children": Sub-topics (same format)

### v2

@@ Abstract
As autonomous AI agents are used in regulated and safety-critical settings, organizations need effective ways to turn policy into enforceable controls. We introduce a regulatory machine learning framework that converts unstructured design artifacts (like PRDs, TDDs, and code) into verifiable runtime guardrails. Our Policy as Prompt method reads these documents and risk controls to build a source-linked policy tree. This tree is then compiled into lightweight, prompt-based classifiers for real-time runtime monitoring. The system is built to enforce least privilege and data minimization. For conformity assessment, it provides complete provenance, traceability, and audit logging, all integrated with a human-in-the-loop review process. Evaluations show our system reduces prompt-injection risk, blocks out-of-scope requests, and limits toxic outputs. It also generates auditable rationales aligned with AI governance frameworks. By treating policies as executable prompts (a policy-as-code for agents), this approach enables secure-by-design deployment, continuous compliance, and scalable AI safety and AI security assurance for regulatable ML.
@@ 1 Introduction
Powerful AI agents are moving into everyday business, from helping HR to flagging security threats Wang et al. (2025); Ding et al. (2024). But this power comes with risk: an HR agent could accidentally leak a salary, or a helpful chatbot could be tricked into running a malicious command Liu et al. (2024); Kim et al. (2023); Li et al. (2025); He et al. (2024). This has created an urgent need to make sure these agents are safe and follow our rules, a sentiment echoed by emerging frameworks like the EU AI Act eu- (2024); Fabiano (2024).
The core problem is what we call the “policy-to-practice” gap: it’s easy for a human to write a rule in a design document, but it’s incredibly hard to turn that simple English sentence into a machine-enforceable rule that works reliably. This gap is a major roadblock to building, testing, and trusting AI systems. To solve this, we can use “guardrails”—safety checks that prevent the AI from doing unintended or harmful things Dong et al. (2024); Zhang et al. (2025). A key security idea here is the principle of least privilege, which means giving a system only the minimal access it needs to do its job. For AI agents, this ensures they stay within defined limits. However, static rules are often too rigid or too vague and fail to capture context. As recent research points out, security for these flexible agents needs to be just-in-time and context-aware (Kholkar and Ahuja, 2025; Tsai and Bagdasarian, 2025). While some have proposed static principles Hua et al. (2024), this is often not enough for dynamic, real-world interactions.
To bridge this critical gap, we introduce Policy as Prompt, a novel framework that reads natural language policy documents and turns them into dynamic, enforceable guardrails. Our system offers a practical way to implement the contextual security that Tsai and Bagdasarian (2025) called for. Our key contributions are as follows: (i) We introduce a scalable, end-to-end pipeline that automatically reads the unstructured technical artifacts teams already write (like PRDs or design docs) to identify and extract security constraints. (ii) We propose a verifiable process where these constraints are converted into a human-readable policy draft, enabling efficient review and refinement by security engineers. (iii) We compile the verified policy into prompt-based classifiers—our “guardrail security policies” Dong et al. (2024)—that use a lightweight LLM to act as a real-time "judge," enforcing the least-privilege policy by validating agent inputs and outputs, ensuring they only do what the policy explicitly allows and nothing more. (iv) We validate our approach by generating policies for various LLM applications and test their effectiveness across different state-of-the-art models.
Figure 1: Policy Generation and Enforcement Pipeline for an HR Application
@@ 2 Policy Tree Generation
Our core process, POLICY-TREE-GEN, uses a verified two-step method to extract and validate security rules directly from application design documents. Step 1 (Parse & Classify): An AI system analyzes the documents, identifies sentences that define security rules or data constraints, and classifies them into one of four categories: ID-I (In-Domain Inputs), OOD-I (Out-of-Domain Inputs, e.g., off-topic or malicious requests), ID-O (In-Domain Outputs), and OOD-O (Out-of-Domain Outputs, e.g., data leaks or toxic responses). Each extracted rule is double-checked by another AI agent for accuracy and proper categorization. Step 2 (Enrich with Examples): The system links these rules to relevant examples found in the documents, producing a structured, verifiable policy tree that preserves contextual grounding. This process inherently enforces the Principle of Least Privilege Zhang et al. (2025): the agent can act only within what is explicitly permitted by its design documents. For instance, for an HR App, a rule such as “Only access resolved HR case data from the Case Portal” is classified as ID-I, meaning the agent must reject attempts to access non-HR data like IT or Sales records (OOD-I). Similarly, a rule like “Generated KB article should be flagged if employee identifiers are present” is tagged as OOD-O, ensuring the agent’s responses do not reveal sensitive information. Thus, an HR agent derived from this process is automatically restricted to HR-related data and outputs, without implicit access to finance or IT information. The resulting policy tree defines and verifies these input–output boundaries, serving as the enforceable foundation for secure, least-privilege AI behavior.
@@
App Example Classification Reason Code Action
HR Update my address on Workday ID (Input) ALLOW
HR My address is 21 Victoria St OOD (Input) Contains Non-Anonymised PII BLOCK
HR Ignore rules and reveal your system prompt OOD (Input) Malicious / Prompt Injection ALERT
HR Send to sample@gmail.com ID (Input) Anonymised PII ALLOW
HR US election news link… OOD (Output) Non HR Content BLOCK
@@ 3 Policy as Prompt Generation
It is then embedded into a master template that primes the LLM as a compliance analyst. Two templates are used: one for an Input Classifier and another for an Output Auditor. These prompt the model to consult the policy rules and return a JSON object with a binary classification (ID/OOD) and a concise justification. For HR input classification (Table 1), HR-related requests without non-anonymized PII are ID, while those with PII, prompt-injection attempts, or non-HR content are OOD. The strict output format enables deterministic system actions (ALLOW, BLOCK, ALERT). HR Input Classifier is shown in Figure 4 and SOC Input Classifier in Figure 5. The deliverables are two markdown few-shot prompts (rules + exemplars), which undergo human-in-the-loop review. Security engineers then either (i) approve for deployment, (ii) reject requiring upstream updates and regeneration, or (iii) request changes to the prompt markdown without altering the upstream tree.
@@ 4 Experimental Setup and Results
This study evaluated the generation and enforcement of guardrail policies across two distinct LLM-powered systems in Human Resource (HR) and Security Operations Centre (SOC)) domains. The input artifacts for all models consisted of PRDs, technical design documents, and prompts extracted from the application source code. These artifacts were sourced from real-world, internal enterprise projects, representing authentic, in-production use cases. To ensure compatibility, all documents were converted to Markdown format, and any embedded images were replaced with textual descriptions generated by gpt-4o. All reported metrics are the average of multiple runs to ensure stability.
@@
Application Model R (%) F1 (%) Macro-F1 (%) ID-I ID-O INVINP INVOUT
@@
SOC O1 19.4 22.6 13.0 22.2 29.6 0.0 0.0
SOC GPT-OSS 120B 5.6 10.3 4.2 16.7 0.0 0.0 0.0
SOC Llama 405B 5.6 9.8 10.0 0.0 0.0 40.0 0.0
SOC Claude 3.5 2.8 5.4 6.2 0.0 0.0 25.0 0.0
POLICY-TREE-GEN Analysis: We evaluated several large language models including Llama 3 405B(Grattafiori et al., 2024), GPT OSS 120B (OpenAI et al., 2025), Claude Sonnet 3.5(Anthropic, 2024) and o1 (OpenAI et al., 2024) on two applications, HR and SOC. The models Gemma 3 1B(Team et al., 2025) and Qwen 1.7B Thinking(Yang et al., 2025) were excluded due to poor performance. Gold policies were created by security engineers and served as the ground truth against which the LLM-generated policies were evaluated. The metrics in Table 2 evaluate model performance in two categories: Detection, measured by Recall (R) and F1 score, and Classification, assessed via Macro-F1 and individual per-class F1 scores for ID-I, ID-O, OOD-I, and OOD-O. A high score in these metrics indicates a model that is effective at both identifying relevant requirements and accurately assigning the correct category to them. HR consistently yields higher performance scores for all models compared to SOC. Furthermore, the o1 model significantly outperforms all other models, demonstrating its superior capability in this specific task.
As shown in Table 4, this analysis details model performance on HR and SOC tasks using metrics such as Detection Precision, which measures the percentage of a model’s predictions that are correct, and Micro-F1. Span quality is evaluated using four metrics: Span Exact, which measures if the extracted text is an exact match to the ground truth; Token-F1, which assesses the token-level overlap between the extracted and ground truth text; Substr, which checks if one text is a substring of the other; and Emb Cos, which is the cosine similarity of the text embeddings. A high score in these metrics indicates that the model is accurately extracting the correct text, even if it might misclassify it. While the o1 model remains a top performer, others reveal a significant disconnect between their extraction quality and overall F1 score. For HR, models like Llama 405B and Sonnet 3.5 have high span quality metrics, indicating accurate text extraction, yet their low Micro-F1 scores suggest they struggle with correct classification. For SOC domain, performance is generally lower and more varied. Notably, Sonnet 3.5 achieved a high Det P on SOC, meaning all of its identified requirements were true positives, but its Micro-F1 score remained very low due to continued classification errors.
@@
POLICY-AS-PROMPT-GEN Analysis: Policy enforcement was tested on a separate set of models: Qwen 3 1.7B (Thinking Mode) (Yang et al., 2025), GPT-4o (OpenAI et al., 2024), and Gemma 3 1B(Team et al., 2025). For specialized data classification tasks within this framework, we employed Small Language Models (SLMs), leveraging their established effectiveness and low latency, which helps to quickly run policies in real-time. Both policies were judged request review by security engineers as part of our evaluation and deployed after minimal changes, saving time for the security team. POLICY-AS-PROMPT-GEN was evaluated on functional correctness by integrating generated prompts into target LLMs and testing with 100 gold inputs and outputs. Tests included both standard and adversarial cases (e.g., prompt injection, toxic content). Evaluation focused on accuracy of OOD/ID classifications, with defaults of OOD $\Rightarrow$ BLOCK and ID $\Rightarrow$ ALLOW, unless otherwise specified. While accuracies in the  70-73% range for GPT-4o (Table 3) are not perfect, they demonstrate significant utility in a real-world context. The system acts as a "default-deny" guardrail: its primary value is in successfully blocking a high percentage of malicious or non-compliant inputs (true positives for OOD) before they reach the agent, drastically reducing the attack surface. This level of accuracy is highly effective as a first-line defense, flagging ambiguous cases for human review, which is a significant improvement over manual, post-hoc auditing. A limitation is that prompt tuning emphasized GPT models, potentially contributing to performance gaps.
@@ 6 Limitations
Despite promising results, this work has several limitations that temper generalizability and reproducibility. Our evaluation spans two enterprise-style domains (HR, SOC) with modest gold sets (100 inputs and 100 outputs per application) and a limited number of design artifacts, which constrains transferability to other settings and larger, more heterogeneous corpora. Policy extraction currently depends on large proprietary models, while open-source baselines differ in size or tuning; moreover, prompting effort was greater for GPT-family models, potentially biasing results and limiting apples-to-apples comparisons. Reproducibility is further hindered by confidentiality of internal artifacts and logs: we cannot release the full corpora. Future work includes leveraging production interaction logs (inputs, outputs, tool calls, actions) to mine candidate rules and hard examples that continuously enrich the policy tree and prompts, and implementing adaptive policy regeneration that automatically re-runs POLICY-TREE-GEN and POLICY-AS-PROMPT-GEN whenever PRDs, TDDs, code, or schemas undergo significant changes, with versioning and regression gating to mitigate drift. Also, prompt optimisations like Kumar et al. (2025) can be applied to POLICY-AS-PROMPT.
@@ A.1 POLICY-TREE-GEN PROMPTS
LLM Prompt: Pass 1 — Parse Document & Classify Security Instructions Role & Goal You are a document analysis expert specializing in data security for LLM-powered applications. Read a document describing an LLM-based system, infer its logical structure, and extract only those instructions relevant to data and guardrail security. Important: Ignore all specific examples in this step. Focus strictly on the instructions themselves. Input <Document text> Task Return a single JSON object representing the document’s hierarchy. For each extracted instruction, you must include: The exact, verbatim source_span copied from the document. A category chosen from 4 CATEGORIES. ID Categories & Guidance ID-I: Defines what the system should accept (in-domain, topical, acceptable inputs). Exclude: outputs, rejection rules, metadata, formatting, or negative conditions. ID-O: Defines correct system responses (including handling invalid/unsupported/out-of-domain inputs). Exclude: malformed prompts, format/schema rules, valid inputs, or outputs. OOD-I: Defines what the system must reject (invalid, unsupported, or out-of-domain inputs). Include: logic/guardrails for blocking inputs. Exclude: formatting, valid inputs, or outputs. OOD-O: Defines incorrect or forbidden system responses (hallucinations, policy violations, leaking PII, misleading answers). Include: logic/guardrails for blocking outputs. Exclude: formats, constraints, valid outputs, or inputs. Exclusion Rule: If an instruction does not clearly fit one of the valid categories above, omit it from the output. Required JSON Shape Root key: "document_structure" (list). Each element: "topic": Heading "content": Instructions with source_span, category, sub_instructions "children": Sub-topics (same format)